OpenStudy (anonymous):

LIMIT x-->-infinite xSin(1/x) as x approaches negative infinity

OpenStudy (ikram002p):

asume x= 1/y y=1/x when x goes to -infinit y goes to 0 LIMIT x-->-infinite xSin(1/x) limit sin y /y as y goes to zero =1

OpenStudy (ikram002p):

so in other words the trick is in changing the limit? yess change the limit

OpenStudy (anonymous):

uhmmm, i had to delete my comment thinking that i have observed wrongly.

OpenStudy (anonymous):

thanx ikram002p. I still want to fint out something from you though, let me close this question.

OpenStudy (ikram002p):

dnt worry i somtimes give wrong answers too , so when u learn sothing there is no wrong question :o, ask wat u dnt got/knew/confused in

OpenStudy (anonymous):

ok, I will not close the question.Tell me what's going on here, if you have an Idea: \[\lim_{\theta \rightarrow 0} Sin5\theta/3\theta \] \[= 5/3 \lim_{\theta \rightarrow 0} ( \sin5\theta/5\theta)\] \[= 1\] \[How?\]

OpenStudy (anonymous):

no , I meant \[= 5/3 \times 1\] \[= 5/3\]

OpenStudy (ikram002p):

ok u have lim sin 5theta/ 3 theta multiply by 5/5 lim 5×sin 5theta/ 5×3 theta 5/3 lim sin 5theta/ 5 theta lim sin 5theta/ 5 theta =1 5/3 lim sin 5theta/ 5 theta =5/3×1

OpenStudy (anonymous):

oh, now I see, thanx again for your help. it was an example given here, and i could not take it in well on some stages. thanx.

OpenStudy (ikram002p):

ur wlc.